A Complete Bibliography of Publications in Chicago Journal of Theoretical Computer Science

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

26 November 2011 Version 1.07

Title word cross-reference

modm [2]. k [12].

-Cut [12].

Abstract [29]. Accuracy [34]. Algorithms [18, 15, 6]. Approximating [12]. Arbitrary [29]. Arithmetic [16]. ATM [7]. Atomicity [9]. Automorphisms [21].

Balancing [35]. Bibliography [20]. Boolean [2]. Bounds [28, 29]. Branching [16].

Channel [10]. Characterizing [33]. Checkable [4]. Checking [37]. Circuits [27, 37, 16]. Classes [33]. Clock [34]. Closed [1]. Coding [31]. Coloring [32]. Combinatorics [15]. Complement [1]. Complements [23]. complete [10]. Completeness [32]. Complexity [37, 25, 24, 15]. Comprehensive [20]. Concurrent [5]. Concurrent-Program [5]. Consistency [29]. Constant [37]. Constant-Depth [37]. Constraint [30]. Context [8]. Context-Sensitive [8]. Correction [13]. Cut [12].

Data [29]. Debate [4]. Decrement [35]. Definition [36]. Depth [33, 37]. Determinant [15]. Distributed [30, 14, 19]. Dual [12]. Dynamic [14].

Efficient [6]. Exclusion [36].

Fair [17]. Finding [21]. Functions [4, 23, 2].

Generating [24]. Grammars [8]. Graph [21, 32]. Graphs [25]. Growing [8].

Hard [4, 6]. Hardness [12]. Heuristics [32]. Higher [33]. Hopfield [26]. Hybrid [29].

Increment [35]. Instances [24]. Isomorphism [16]. Isomorphisms [22].

Large [27]. Layout [7]. Lemma [31]. Limited [11]. Linear [28]. Local [36]. Logspace [1].

Manhattan [10]. Max [12]. Measures [3, 5]. Multitolerance [19]. Multivalued [23]. Mutual [36].

Network [18]. Networks [35, 7, 9, 26]. Neural [26]. Nonapproximability [4]. Nondeterminism [11]. NP [10]. NP-complete [10]. Number [21].

OBDDs [25]. Once [16]. Operations [35]. Operators [33]. Optimal [7]. Orientation [9]. Orthogonal [34].

P [6]. Paper [20]. Path [7]. Permanent [27]. Polynomial [11]. Power [18]. Predicates [5]. Probabilistically [4]. Problem [16]. Problems [28, 25]. Program [37, 5]. Programs [16]. Progress [5]. Protocols [14]. PSPACE [4]. PSPACE-Hard [4].

Query [37].

Rabin [3]. Randomized [22]. Rank [5]. Read [9, 16]. Read-Once [16]. Read/Write [9]. Reductions [22]. Refinement [36]. Representation [2].

Represented [25]. Requires [27]. Reset [19]. Restricted [10]. Routing [7, 10].

Satisfaction [30]. Satisfiability [28, 31]. Self [18, 36, 30, 20, 9, 26, 13]. Self-Stabilization [20, 26, 13]. Self-Stabilizing [18, 36, 30, 9]. Sensitive [8]. Sets [6]. Settings [10]. Shared [7]. Small [33]. Space [33, 6]. Space-Efficient [6]. Sparse [6]. Stabilization [20, 26, 13]. Stabilizing [18, 36, 30, 9]. Strong [29]. Superstabilizing [14]. Supply [18]. Supporting [35]. Switches [7]. Symmetric [1]. Synchronization [34]. Systems [4, 14, 17].

Table [7]. Test [24]. Threshold [27]. Time [29]. Transition [17]. Tree [13]. Truly [10]. Type [33]. Types [29].

Unicyclic [9]. Unidirectional [18].

Uniform [27, 9].

Verification [17, 5]. Versus [32, 11].

Virtual [7]. vs [5].

Weak [2]. Weakly [8]. Working [20].

Write [9].

Yield [6].

References

Nisan:1995:SIC

[1] Noam Nisan and Amnon Ta-Shma. Symmetric logspace is closed under complement. Chicago Journal of Theoretical Computer Science, June 30, 1995. ISSN 1073-0486. URL http://www.cs.

                                            uchicago.edu/publications/cjtcs/
                                            articles/1995/1/contents.html.

Grolmusz:1995:WMI

[2] Vince Grolmusz. On the weak mod m representation of Boolean functions. Chicago Journal of Theoretical Computer Science, July 21, 1995. ISSN 1073-0486.

URL http://www.cs.uchicago.edu/
publications/cjtcs/articles/1995/
2/contents.html.

Klarlund:1995:RM

[3] Nils Klarlund and Dexter Kozen. Rabin measures. Chicago Journal of Theoretical Computer Science, September 20, 1995. ISSN 1073-0486.

URL http://www.cs.uchicago.edu/
publications/cjtcs/articles/1995/
3/contents.html.

Condon:1995:PCD

[4] Anne Condon, Joan Feigenbaum, Carsten Lund, and Peter W. Shor. Probabilistically checkable debate systems and nonapproximability of PSPACEhard functions. Chicago Journal of Theoretical Computer Science, October 19, 1995. ISSN 1073-0486.

URL http://www.cs.uchicago.edu/
publications/cjtcs/articles/1995/
4/contents.html.

Vardi:1996:RPV

[5] Moshe Y. Vardi. Rank predicates vs. progress measures in concurrentprogram verification. Chicago Journal of Theoretical Computer Science, February 9, 1996. ISSN 1073-0486.

URL http://www.cs.uchicago.edu/
publications/cjtcs/articles/1996/
1/contents.html.

Ogihara:1996:SHS

[6] Mitsunori Ogihara. Sparse hard sets for P yield space-efficient algorithms. Chicago Journal of Theoretical Computer Science, March 27, 1996. ISSN 1073-0486.

URL http://www.cs.uchicago.edu/
publications/cjtcs/articles/1996/
2/contents.html.

Gerstel:1996:OVP

[7] Ornan Gerstel, Israel Cidon, and Shmuel Zaks. Optimal virtual path layout in ATM networks with shared routing table switches. Chicago Journal of Theoretical Computer Science, October 31, 1996. ISSN 1073-0486.

URL http://www.cs.uchicago.edu/
publications/cjtcs/articles/1996/

3/contents.html. Published in Selected Papers from PODC 1994, David Peleg editor.

Buntrock:1996:WGC

[8] Gerhard Buntrock and Gundula Niemann. Weakly growing contextsensitive grammars. Chicago Journal of Theoretical Computer Science, November 13, 1996. ISSN 1073-0486.

URL http://www.cs.uchicago.edu/
publications/cjtcs/articles/1996/
4/contents.html.

Hoover:1996:USS

[9] H. James Hoover and Piotr Rudnicki. Uniform self-stabilizing orientation of unicyclic networks under read/write atomicity. Chicago Journal of Theoretical Computer Science, December 5, 1996. ISSN 1073-0486.

URL http://www.cs.uchicago.edu/
publications/cjtcs/articles/1996/

5/contents.html. Special Issue on Self-Stabilization, Shlomi Dolev and Jennifer Welch editors.

Middendorf:1996:MCR

[10] Martin Middendorf. Manhattan channel routing is NP-complete under truly restricted settings. Chicago Journal of Theoretical Computer Science, December 30, 1996. ISSN 1073-0486.

URL http://www.cs.uchicago.edu/
publications/cjtcs/articles/1996/
6/contents.html.

Feige:1997:LVP

[11] Uriel Feige and Joe Kilian. On limited versus polynomial nondeterminism. Chicago Journal of Theoretical Computer Science, March 12, 1997. ISSN 1073-0486.

URL http://www.cs.uchicago.edu/
publications/cjtcs/articles/1997/
1/contents.html.

Kann:1997:HAM

[12] Viggo Kann, Sanjeev Khanna, Jens Lagergren, and Alessandro Panconesi. On the hardness of approximating max k-cut and its dual. Chicago Journal of Theoretical Computer Science, June 3, 1997. ISSN 1073-0486.

URL http://www.cs.uchicago.edu/
publications/cjtcs/articles/1997/
2/contents.html.

Varghese:1997:SST

[13] George Varghese, Anish Arora, and Mohamed Gouda. Self-stabilization by tree correction. Chicago Journal of Theoretical Computer Science, November 4, 1997. ISSN 1073-0486.

URL http://www.cs.uchicago.edu/
publications/cjtcs/articles/1997/

3/contents.html. Special Issue on Self-Stabilization, Shlomi Dolev and Jennifer Welch editors.

Dolev:1997:SPD

[14] Shlomi Dolev and Ted Herman. Superstabilizing protocols for dynamic distributed systems. Chicago Journal of Theoretical Computer Science, December 19, 1997. ISSN 1073-0486.

URL http://www.cs.uchicago.edu/
publications/cjtcs/articles/1997/

4/contents.html. Special Issue on Self-Stabilization, Shlomi Dolev and Jennifer Welch editors.

Mahajan:1997:DCA

[15] Meena Mahajan and V. Vinay. Determinant: Combinatorics, algorithms, and complexity. Chicago Journal of Theoretical Computer Science, December 31, 1997. ISSN 1073-0486.

URL http://www.cs.uchicago.edu/
publications/cjtcs/articles/1997/
5/contents.html.

Thierauf:1998:IPR

[16] Thomas Thierauf. The isomorphism problem for read-once branching programs and arithmetic circuits. Chicago Journal of Theoretical Computer Science, May 16, 1998. ISSN 1073-0486.

URL http://www.cs.uchicago.edu/
publications/cjtcs/articles/1998/

1/contents.html. Special Issue on Computational Complexity, from the 1996 Dagstuhl-Seminar, Eric Allender editor.

Kupferman:1998:VFT

[17] Orna Kupferman and Moshe Y. Vardi. Verification of fair transition systems. Chicago Journal of Theoretical Computer Science, March 16, 1998. ISSN 1073-0486.

URL http://www.cs.uchicago.edu/
publications/cjtcs/articles/1998/
2/contents.html.

Afek:1998:SSU

[18] Yehuda Afek and Anat Bremler. Selfstabilizing unidirectional network algorithms by power supply. Chicago Journal of Theoretical Computer Science, December 7, 1998. ISSN 1073-0486.

URL http://www.cs.uchicago.edu/
publications/cjtcs/articles/1998/

3/contents.html. Special Issue on Self-Stabilization, Shlomi Dolev and Jennifer Welch editors.

Kulkarni:1998:MDR

[19] Sandeep S. Kulkarni and Anish Arora.

Multitolerance in distributed reset. Chicago Journal of Theoretical Computer Science, December 7, 1998. ISSN 1073-

0486. URL http://www.cs.uchicago.
edu/publications/cjtcs/articles/
1998/4/contents.html. Special Issue on

Self-Stabilization, Shlomi Dolev and Jennifer Welch editors.

Herman:1998:WPC

[20] Ted Herman. Working paper: A comprehensive bibliography on self-stabilization. Chicago Journal of Theoretical Computer Science, 1998. ISSN 1073-

0486. URL http://www.cs.uiowa.edu/
ftp/selfstab/bibliography/.

Beals:1999:FNG

[21] Robert Beals, Richard Chang, William Gasarch, and Jacobo Torán. On finding the number of graph automorphisms. Chicago Journal of Theoretical Computer Science, February 10, 1999. ISSN 1073-

0486. URL http://www.cs.uchicago.
edu/publications/cjtcs/articles/
1999/1/contents.html.

Wang:1999:RRI

[22] Jie Wang. Randomized reductions and isomorphisms. Chicago Journal of Theoretical Computer Science, February 24, 1999. ISSN 1073-0486.

URL http://www.cs.uchicago.edu/
publications/cjtcs/articles/1999/
2/contents.html. Special Issue on

Computational Complexity, results from Dagstuhl-Seminar 1996, Eric Allender editor.

Fenner:1999:CMF

[23] Stephen Fenner, Frederic Green, Steven Homer, Alan L. Selman, Thomas Thierauf, and Heribert Vollmer. Complements of multivalued functions. Chicago Journal of Theoretical Computer Science, March 19, 1999. ISSN 1073-0486.

URL http://www.cs.uchicago.edu/
publications/cjtcs/articles/1999/
3/contents.html.

Karg:1999:CGT

[24] Christoph Karg, Johannes Köbler, and Rainer Schuler. The complexity of generating test instances. Chicago Journal of Theoretical Computer Science, April 22, 1999. ISSN 1073-0486.

URL http://www.cs.uchicago.edu/
publications/cjtcs/articles/1999/
4/contents.html. Special Issue on

Computational Complexity, results from Dagstuhl-Seminar 1996, Eric Allender editor.

Feigenbaum:1999:CPG

[25] Joan Feigenbaum, Sampath Kannan, Moshe Y. Vardi, and Mahesh Viswanathan. The complexity of problems on graphs represented as OBDDs. Chicago Journal of Theoretical Computer Science, August 6, 1999. ISSN 1073-0486.

URL http://www.cs.uchicago.edu/
publications/cjtcs/articles/1999/
5/contents.html.

Jagota:1999:HNN

[26] Arun Jagota. Hopfield neural networks and self-stabilization. Chicago Journal of Theoretical Computer Science, August 6, 1999. ISSN 1073-0486.

URL http://www.cs.uchicago.edu/
publications/cjtcs/articles/1999/
6/contents.html.

Allender:1999:PRL

[27] Eric Allender. The permanent requires large uniform threshold circuits. Chicago Journal of Theoretical Computer Science, August 6, 1999. ISSN 1073-0486.

URL http://www.cs.uchicago.edu/
publications/cjtcs/articles/1999/
7/contents.html.

Erickson:1999:BLS

[28] Jeff Erickson. Bounds for linear satisfiability problems. Chicago Journal of Theoretical Computer Science, August 6, 1999. ISSN 1073-0486.

URL http://www.cs.uchicago.edu/
publications/cjtcs/articles/1999/
8/contents.html.

Kosa:1999:TBS

[29] Martha J. Kosa. Time bounds for strong and hybrid consistency for arbitrary abstract data types. Chicago Journal of Theoretical Computer Science, August 6, 1999. ISSN 1073-0486.

URL http://www.cs.uchicago.edu/
publications/cjtcs/articles/1999/
9/contents.html.

Collin:1999:SSD

[30] Zeev Collin, Rina Dechter, and Shmuel Katz. Self-stabilizing distributed constraint satisfaction. Chicago Journal of Theoretical Computer Science, December 31, 1999. ISSN 1073-0486.

URL http://www.cs.uchicago.edu/
publications/cjtcs/articles/1999/
10/contents.html.

Paturi:1999:SCL

[31] Ramamohan Paturi, Pavel Pudlak, and Francis Zane. Satisfiability coding lemma. Chicago Journal of Theoretical Computer Science, December 31, 1999. ISSN 1073-

0486. URL http://www.cs.uchicago.
edu/publications/cjtcs/articles/
1999/11/contents.html.

Rothe:2000:HVC

[32] Jorg Rothe. Heuristics versus completeness for graph coloring. Chicago Journal of Theoretical Computer Science, February 29, 2000. ISSN 1073-

0486. URL http://www.cs.uchicago.
edu/research/publications/cjtcs/
articles/2000/1/contents.html.

Agrawal:2000:CSD

[33] Manindra Agrawal, Eric Allender, Samir Datta, Heribert Vollmer, and Klaus W. Wagner. Characterizing small depth and small space classes by operators of higher type. Chicago Journal of Theoretical Computer Science, September 6, 2000. ISSN 1073-

0486. URL http://www.cs.uchicago.
edu/research/publications/cjtcs/
articles/2000/2/contents.html.

Schmid:2000:OAC

[34] Ulrich Schmid. Orthogonal accuracy clock synchronization. Chicago Journal of Theoretical Computer Science, August 17, 2000. ISSN 1073-

0486. URL http://www.cs.uchicago.

edu/research/publications/cjtcs/
articles/2000/3/contents.html.

Aiello:2000:SID

[35] William Aiello, Costas Busch, Maurice Herlihy, Marios Mavronicolas, Nir Shavit, and Dan Touitou. Supporting increment and decrement operations in balancing networks. Chicago Journal of Theoretical Computer Science, December 14, 2000. ISSN 1073-

0486. URL http://www.cs.uchicago.
edu/research/publications/cjtcs/
articles/2000/4/contents.html.

Beauquier:2002:SSL

[36] Jeffrey Beauquier, Ajoy K. Datta, Maria Gradinariu, and Frederic Magniette. Self-stabilizing local mutual exclusion and definition refinement. Chicago Journal of Theoretical Computer Science, July 24, 2002. ISSN 1073-0486.

URL http://cjtcs.cs.uchicago.edu/
articles/2002/1/contents.html.

Arvind:2002:QCP

[37] V. Arvind, K. V. Subrahmanyam, and N. V. Vinodchandran. The query complexity of program checking by constantdepth circuits. Chicago Journal of Theoretical Computer Science, December 5, 2002. ISSN 1073-0486. URL http:

//cjtcs.cs.uchicago.edu/articles/
2002/2/contents.html.